OpenStudy (anonymous):

A car traveling at 6.7m/s accelerates 2.7m/s^2 to reach a speed of 11.0m/s. How long does it take for this to acceleration to occur?

OpenStudy (anonymous):

It's another problem based around the definition of acceleration

OpenStudy (anonymous):

Yes but this one I believe uses V=v1+at or something like that

OpenStudy (anonymous):

well, maybe, but it's the same thing really

OpenStudy (anonymous):

you want to use Vfinal = Vinitial + at, right ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

but I can rearrange your equation as a = (Vfinal-Vinitial)/t, which is just the definition of acceleration : )

OpenStudy (anonymous):

However, in your question, we know a and we know Vfinal and Vinitial, so it's just a matter of rearranging the equation so that we can plug in the numbers and figure out t

OpenStudy (anonymous):

11.0= 6.7-2.7/ t

OpenStudy (anonymous):

no, that doesn't look right

OpenStudy (anonymous):

6.7= 11.0-2.7/ t?

OpenStudy (anonymous):

If you rearrange things to find t you should have t = (Vfinal-Vinitial)/a It is important to put the brackets in

OpenStudy (anonymous):

T= 6.7-2.7/11.0?

OpenStudy (anonymous):

what is Vfinal ?

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

Well it is given pretty clearly in the question

OpenStudy (anonymous):

11.0?

OpenStudy (anonymous):

yes, Vfinal is 11 so what about Vinitial ?

OpenStudy (anonymous):

2.7

OpenStudy (anonymous):

are you just guessing or have you reread the question ?

OpenStudy (anonymous):

I'm guessing because I wasn't there the day we learned this so I have no idea what any of this means

OpenStudy (anonymous):

well that's no way to tackle problems in physics, you need to have some idea what it is all about first but in any case, you first need to carefully read the question and absorb what it is telling you

OpenStudy (anonymous):

Okay, would it be t= (11.0-6.7)/2.7

OpenStudy (anonymous):

yes, that is right I'm sorry you're having trouble with this, hope it becomes clearer for you with practice try to keep a cool head

OpenStudy (anonymous):

Okay so then I got 1.59s and that makes sense I think

OpenStudy (anonymous):

yes you've got the right numbers